OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: y ≤ –x + 7 y < 3x + 11 A (5, 5) B (–5, 5) C (–5, –5) D none of the above

OpenStudy (angela210793):

5<=-5+7 5<15+11 5<=2 not true 5<26 true not a solution 5<=5+7 5<-15+11 5<=12 true 5<-4 not true not a solution -5<=5+7 -5<-15+11 -5<=12 true -5<-4 true C(-5;-5) is a solution :D